b. What is the mass in grams of H2 when 4.0 moles of HCl is added to the reaction?

1
Expert's answer
2021-12-15T10:03:48-0500

The reaction is:

Mg(s) + 2HCl(aq) = MgCl2(aq) + H2(g)

The chemical equation shows that: 1 mole of Mg + 2 moles of HCl = 1 mole of H2 + 1 mole of MgCl2


Moles of "H_2=\frac{4.0}{2}=2.0moles"


The molar mass of hydrogen gas is 2.016 grams per mole


Mass"=2.016\u00d72=4.032grams"